# On Minimax Estimation of Parameters in Softmax-Contaminated Mixture of Experts

### v1

@@ A.3 Proof of Theorem 3
$[\exp(\tau_{n}){(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(u)}(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(v)}}+[\exp(\tau^{*}_{n})-\exp(\tau_{n})](-\Delta\eta^{*}_{n})^{(u)}(-\Delta\eta^{*}_{n})^{(v)}]/m_{n}$ $\to\beta_{10,0uv},$
@@
$[\exp(\tau_{n})(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(u)}(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(v)}+[\exp(\tau^{*}_{n})-\exp(\tau_{n})](-\Delta\eta^{*}_{n})^{(u)}(-\Delta\eta^{*}_{n})^{(v)}]/m_{n}$ $\to\beta_{20,0uv},$
$[\exp(\tau_{n})(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(u)}(\Delta\nu_{n}-\Delta\nu^{*}_{n})+[\exp(\tau^{*}_{n})-\exp(\tau_{n})](-\Delta\eta^{*}_{n})^{(u)}(-\Delta\nu^{*}_{n})]/m_{n}$ $\to\beta_{30,0u0},$
@@
$0=\lim_{n\to\infty}\frac{1}{m_{n}}\frac{2\mathbb{E}_{X}[V(p_{G_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]}{{{D_{2}}(G_{n},G_{*,n})}}\geq\int\liminf_{n\to\infty}\frac{1}{m_{n}}\frac{|p_{G_{n}}(Y|X)-p_{G_{*,n}}(Y|X)|}{{{D_{2}}(G_{n},G_{*,n})}}d(X,Y).$
@@
$\mathbb{E}_{X}[V(p_{G_{n}}(\cdot|X),p_{G_{*,n}}(\cdot|X))]/D_{2}(G_{n},G_{n}^{*})\to 0,$
@@
$D_{2}(G_{n},G_{n}^{*})\to D_{2}(G,G_{*})\neq 0,\mathbb{E}_{X}[V(p_{G_{n}}(\cdot|X),p_{G_{*,n}}(\cdot|X))]\to\mathbb{E}_{X}[V(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))],$
@@
$f(Y|h(X,\eta_{0}),\nu_{0})=\dfrac{1}{1+\exp(\beta^{\top}X+\tau^{*})}f(Y|h(X,\eta_{0}),\nu_{0})+\dfrac{\exp(\beta^{\top}X+\tau^{*})}{1+\exp(\beta^{\top}X+\tau^{*})}f(Y|h(X,\eta),\nu)$
$\Rightarrow$ $\dfrac{\exp(\beta^{\top}X+\tau^{*})}{1+\exp(\beta^{\top}X+\tau^{*})}f(Y|h(X,\eta_{0}),\nu_{0})=\dfrac{\exp(\beta^{\top}X+\tau^{*})}{1+\exp(\beta^{\top}X+\tau^{*})}f(Y|h(X,\eta),\nu)$
$\Rightarrow$ $f(Y|h(X,\eta_{0}),\nu_{0})=f(Y|h(X,\eta),\nu)\text{ (as $\exp(\beta^{\top}X+\tau^{*})\neq 0$)}$
@@ A.4 Proof of Theorem 4
In what follows, we present a proof of Theorem 4 for the non-distinguishable setting.
@@
    such that $d_{\prime}(S_{1,n},S_{2,n})\to 0$ and $\mathbb{E}_{X}[h(p_{S_{1,n}}(\cdot|X),p_{S_{2,n}}(\cdot|X))]/d_{\prime}^{r}(S_{1,n},S_{2,n})\to 0$ as $n\to\infty$.
@@
    such that $d_{\prime\prime}(S_{1,n},S_{2,n})\to 0$ and $\mathbb{E}_{X}[h(p_{S_{1,n}}(\cdot|X),p_{S_{2,n}}(\cdot|X))]/d_{\prime\prime}^{r}(S_{1,n},S_{2,n})\to 0$ as $n\to\infty$.
@@ B.1 Proof of Proposition 1
The ratio $g_{1}/g_{2}$ is the analytic (non-polynomial) function
@@
[…]
@@ B.2 Proof of Proposition 2
$p_{G}(y|x)=\bigl{[}1-\lambda(x)\bigr{]}\,f_{0}\!\bigl{(}y|h_{0}(x,\eta_{0}),\nu_{0}\bigr{)}+\lambda(x)\,f\!\bigl{(}y|h(x,\eta),\nu\bigr{)},\qquad\lambda(x):=\frac{\exp\!\bigl{(}\beta^{\top}x+\tau\bigr{)}}{1+\exp\!\bigl{(}\beta^{\top}x+\tau\bigr{)}},$
$p_{G^{\prime}}(y|x)=\bigl{[}1-\lambda^{\prime}(x)\bigr{]}\,f_{0}\!\bigl{(}y|h_{0}(x,\eta_{0}),\nu_{0}\bigr{)}+\lambda^{\prime}(x)\,f\!\bigl{(}y|h(x,\eta^{\prime}),\nu^{\prime}\bigr{)},\qquad\lambda^{\prime}(x):=\frac{\exp\!\bigl{(}{\beta^{\prime}}^{\top}x+\tau^{\prime}\bigr{)}}{1+\exp\!\bigl{(}{\beta^{\prime}}^{\top}x+\tau^{\prime}\bigr{)}}.$
@@
$f_{0}\!\bigl{(}y|h_{0},\nu_{0}\bigr{)}+b(x)\,f\!\bigl{(}y|h(x,\eta^{\prime}),\nu^{\prime}\bigr{)}+c(x)\,f\!\bigl{(}y|h(x,\eta),\nu\bigr{)}=0,\quad b(x):=\frac{\lambda^{\prime}(x)}{\lambda^{\prime}(x)-\lambda(x)},\;c(x):=\frac{-\lambda(x)}{\lambda^{\prime}(x)-\lambda(x)}.$
@@
Because the soft-max map $(\beta,\tau)\mapsto\lambda(\cdot)$ is injective, we conclude
@@ B.3 Proof of Proposition 3
[…]
@@
$\sup_{G_{*}\in\Xi}\mathbb{E}_{p_{G_{*},n}}\mathbb{E}_{X}[h(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\leq C\sqrt{\log n/n}.$
@@
$H_{B}(\epsilon,\mathcal{P}(\Xi),h)\lesssim\log(1/\epsilon),$ (55)
@@
    $H_{B}(\epsilon,\mathcal{P}(\Xi),h)\lesssim\log(1/\epsilon)$.
@@
Now we will prove Lemma 3, Lemma 4 and Lemma 5 in order. At first we need to introduce another Lemma 6 before we prove Lemma 3. Lemma 6 is Theorem 5.11 in Vandegeer-2000 and its proof can also be found in Vandegeer-2000 .
@@
Let $R>0$, $k\geq 1$ and $\mathcal{G}$ is a subset in $\Xi$ where $G_{*}\in\mathcal{G}\subset\Xi$ Given $C_{1}<\infty$, for all $C$ sufficiently large, and for $n\in\mathbb{N}$ and $t>0$ is in the following range
@@
[…]
@@
Firstly, by Lemma 4.1 and 4.2 in Vandegeer-2000 , we have
$\frac{1}{16}\mathbb{E}_{X}[h^{2}(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\leq\mathbb{E}_{X}[h^{2}(\bar{p}_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\leq\frac{1}{\sqrt{n}}\mu_{n}(\widehat{G}_{n}),$
@@
$\mathbb{P}_{G_{*,n}}(\mathbb{E}_{X}[h(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\geq\delta)$
$\leq\mathbb{P}_{G_{*,n}}\left(\mu_{n}(\widehat{G}_{n})-\sqrt{n}\mathbb{E}_{X}[h^{2}(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\geq 0,\mathbb{E}_{X}[h(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\geq\frac{\delta}{4}\right)$
$\leq\mathbb{P}_{G_{*,n}}\left(\sup_{G:\mathbb{E}_{X}[h(\bar{p}_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]\geq\delta/4}\left[\mu_{n}(G)-\sqrt{n}\mathbb{E}_{X}[h^{2}(\bar{p}_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]\right]\geq 0\right)$
$\leq\sum_{s=0}^{S}\mathbb{P}_{G_{*,n}}\left(\sup_{G:2^{s}\delta/4\leq\mathbb{E}_{X}[h(\bar{p}_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]\leq 2^{s+1}\delta/4}\left|\mu_{n}(G)\right|\geq\sqrt{n}2^{2s}(\frac{\delta}{4})^{2}\right)$
$\leq\sum_{s=0}^{S}\mathbb{P}_{G_{*,n}}\left(\sup_{G:\mathbb{E}_{X}[h(\bar{p}_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]\leq 2^{s+1}\delta/4}\left|\mu_{n}(G)\right|\geq\sqrt{n}2^{2s}(\frac{\delta}{4})^{2}\right)$
where $S$ is a smallest number such that $12^{S}\delta/4>$ 1 .
@@
[…]
@@
$\mathbb{P}_{G_{*,n}}\left(\mathbb{E}_{X}[h(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]>\delta\right)\leq C\sum_{s=0}^{\infty}\exp\left(-\frac{2^{2s}n\delta^{2}}{2^{14}C^{2}}\right)\leq c\exp\left(-\frac{n\delta^{2}}{c}\right),$ (59)
@@
$\mathbb{E}_{p_{G_{*,n}}}\mathbb{E}_{X}[h({p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X)})]$ $=\int^{\infty}_{0}\mathbb{P}\left(\mathbb{E}_{X}[h({p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X)})]>\delta\right)d\delta$
[…]
@@
$\sup_{G_{*}\in\Xi}\mathbb{E}_{p_{G_{*},n}}\mathbb{E}_{X}[h(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\leq C\sqrt{\log n/n}.$
@@
Now, using the fact that for densities $f^{*},f_{1},f_{2}$, we have $h^{2}\left(\frac{f_{1}+f^{*}}{2},\frac{f_{2}+f^{*}}{2}\right)\leq\frac{h^{2}(f_{1},f_{2})}{2}$, it is easy to verify that $H_{B}(\delta/\sqrt{2},\overline{\mathcal{P}}(\Xi),h)\leq H_{B}(\delta,\mathcal{P}(\Xi),h)$. Hence, if equation (55) holds true, then
$H_{B}(\delta,\overline{\mathcal{P}}^{1/2}(\Xi,\delta),\mu)\leq H_{B}(\delta,{\mathcal{P}}(\Xi),h)\lesssim\log\left(\frac{1}{\delta}\right).$
@@
Let $\mathcal{P}(\Theta):=\{p_{\Upsilon}:\Upsilon\in\Theta\}$, where $p_{\Upsilon}(Y|X):=f(Y|h(X,\eta),\nu)$. By Lemma 6 in ho2022gaussian , we have
@@
$\quad+\sigma_{\widetilde{\lambda}}(X)\left[f(Y|h(X,\eta),\nu)-f(Y|h(X,\widetilde{\eta}),\widetilde{\nu})\right],$
@@
here $C_{1}$ is a positive constant depending on $\ell$ and $f_{0}$. Moreover $H(X,Y)$ is an envelope of $\mathcal{P}(\Xi)$. Next, let $g_{1},\dots,g_{N}$ represent an $\eta$-net over $\mathcal{P}_{k}(\Xi)$. Then, we construct the brackets $[p^{L}_{i}(X,Y),p^{U}_{i}(X,Y)]$ as follows:

### v2

@@ A.3 Proof of Theorem 3
$[\exp(\tau_{n}){(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(u)}(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(v)}}+[\exp(\tau^{*}_{n})-\exp(\tau_{n})](-\Delta\eta^{*}_{n})^{(u)}(-\Delta\eta^{*}_{n})^{(v)}]/m_{n}$
$\to\beta_{10,0uv},$
@@
$[\exp(\tau_{n})(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(u)}(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(v)}+[\exp(\tau^{*}_{n})-\exp(\tau_{n})](-\Delta\eta^{*}_{n})^{(u)}(-\Delta\eta^{*}_{n})^{(v)}]/m_{n}$
$\to\beta_{20,0uv},$
$[\exp(\tau_{n})(\Delta\eta_{n}-\Delta\eta^{*}_{n})^{(u)}(\Delta\nu_{n}-\Delta\nu^{*}_{n})+[\exp(\tau^{*}_{n})-\exp(\tau_{n})](-\Delta\eta^{*}_{n})^{(u)}(-\Delta\nu^{*}_{n})]/m_{n}$
$\to\beta_{30,0u0},$
@@
$0=\lim_{n\to\infty}\frac{1}{m_{n}}\frac{2\mathbb{E}_{X}[d_{V}(p_{G_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]}{{{D_{2}}(G_{n},G_{*,n})}}\geq\int\liminf_{n\to\infty}\frac{1}{m_{n}}\frac{|p_{G_{n}}(Y|X)-p_{G_{*,n}}(Y|X)|}{{{D_{2}}(G_{n},G_{*,n})}}d(X,Y).$
@@
$\mathbb{E}_{X}[h_{V}(p_{G_{n}}(\cdot|X),p_{G_{*,n}}(\cdot|X))]/D_{2}(G_{n},G_{n}^{*})\to 0,$
@@
$D_{2}(G_{n},G_{n}^{*})\to D_{2}(G,G_{*})\neq 0,\mathbb{E}_{X}[h_{V}(p_{G_{n}}(\cdot|X),p_{G_{*,n}}(\cdot|X))]\to\mathbb{E}_{X}[h_{V}(p_{G}(\cdot|X),p_{G_{*}}(\cdot|X))],$
@@
$f(Y|h(X,\eta_{0}),\nu_{0})$ $=\frac{1}{1+\exp(\beta^{\top}X+\tau^{*})}f(Y|h(X,\eta_{0}),\nu_{0})+\frac{\exp(\beta^{\top}X+\tau^{*})}{1+\exp(\beta^{\top}X+\tau^{*})}f(Y|h(X,\eta),\nu)$
which further implies that
$\frac{\exp(\beta^{\top}X+\tau^{*})}{1+\exp(\beta^{\top}X+\tau^{*})}f(Y|h(X,\eta_{0}),\nu_{0})$ $=\frac{\exp(\beta^{\top}X+\tau^{*})}{1+\exp(\beta^{\top}X+\tau^{*})}f(Y|h(X,\eta),\nu)$
and hence
$f(Y|h(X,\eta_{0}),\nu_{0})$ $=f(Y|h(X,\eta),\nu)\quad\text{(as $\exp(\beta^{\top}X+\tau^{*})\neq 0$)}.$
@@ A.4 Proof of Theorem 4
In what follows, we present the proof of Theorem 4 for the non-distinguishable setting.
@@
    such that $d_{\prime}(S_{1,n},S_{2,n})\to 0$ and $\mathbb{E}_{X}[h_{H}(p_{S_{1,n}}(\cdot|X),p_{S_{2,n}}(\cdot|X))]/d_{\prime}^{r}(S_{1,n},S_{2,n})\to 0$ as $n\to\infty$.
@@
    such that $d_{\prime\prime}(S_{1,n},S_{2,n})\to 0$ and $\mathbb{E}_{X}[h_{H}(p_{S_{1,n}}(\cdot|X),p_{S_{2,n}}(\cdot|X))]/d_{\prime\prime}^{r}(S_{1,n},S_{2,n})\to 0$ as $n\to\infty$.
@@ B.1 Proof of Proposition 1
The ratio $g_{1}/g_{2}$ is the analytic (non‑polynomial) function
@@
with $K\neq 0$. Since $\mu_{1}\neq\mu_{2}$ or $\sigma_{1}^{2}\neq\sigma_{2}^{2}$, this exponential term cannot be expressed as a quadratic polynomial in $y$. Consequently the set of functions $\bigl\{g_{1}/g_{2},1,y-\mu_{2},(y-\mu_{2})^{2}\bigr\}$ is linearly independent on any interval. Hence every coefficient in the polynomial identity must vanish:
@@ B.2 Proof of Proposition 2
$p_{G}(y|x)$ $=\bigl[1-\lambda(x)\bigr]f_{0}\bigl(y|h_{0}(x,\eta_{0}),\nu_{0}\bigr)+\lambda(x)f\bigl(y|h(x,\eta),\nu\bigr),$
$p_{G^{\prime}}(y|x)$ $=\bigl[1-\lambda^{\prime}(x)\bigr]f_{0}\bigl(y|h_{0}(x,\eta_{0}),\nu_{0}\bigr)+\lambda^{\prime}(x)f\bigl(y|h(x,\eta^{\prime}),\nu^{\prime}\bigr),$
where $\lambda(x):=\frac{\exp\bigl(\beta^{\top}x+\tau\bigr)}{1+\exp\bigl(\beta^{\top}x+\tau\bigr)}$ and $\lambda^{\prime}(x):=\frac{\exp\bigl({\beta^{\prime}}^{\top}x+\tau^{\prime}\bigr)}{1+\exp\bigl({\beta^{\prime}}^{\top}x+\tau^{\prime}\bigr)}$.
@@
$f_{0}\bigl(y|h_{0},\nu_{0}\bigr)+b(x)f\bigl(y|h(x,\eta^{\prime}),\nu^{\prime}\bigr)+c(x)f\bigl(y|h(x,\eta),\nu\bigr)=0,$
where
$b(x):=\frac{\lambda^{\prime}(x)}{\lambda^{\prime}(x)-\lambda(x)},c(x):=\frac{-\lambda(x)}{\lambda^{\prime}(x)-\lambda(x)}.$
@@
Because the soft‑max map $(\beta,\tau)\mapsto\lambda(\cdot)$ is injective, we conclude
@@ B.3 Proof of Proposition 3
In particular, let $\mathcal{P}(\Xi):=\{p_{\lambda}:\lambda\in\Xi\}$, and define the symmetrized density $\bar{p}_{\lambda}:=\frac{1}{2}(p^{*}+p_{\lambda}),$ where $p^{*}$ denotes the true density. We then define the following sets: $\overline{\mathcal{P}}(\Xi):=\{\bar{p}_{\lambda}:\lambda\in\Xi\}$ and $\overline{\mathcal{P}}^{1/2}(\Xi):=\{\bar{p}_{\lambda}^{1/2}:\bar{p}_{\lambda}\in\overline{\mathcal{P}}(\Xi)\}.$ To study convergence rates, we consider the localized version of the symmetrized class: $\overline{\mathcal{P}}^{1/2}(\Xi,\epsilon):=\{\bar{p}_{\lambda}^{1/2}\in\overline{\mathcal{P}}^{1/2}(\Xi):d_{H}(\bar{p}_{\lambda},p^{*})\leq\epsilon\},$ where $d_{H}(\cdot,\cdot)$ denotes the Hellinger distance. Then we assess the complexity of this class via the bracketing entropy integral defined in [Vandegeer-2000]: $\mathcal{J}_{B}(\epsilon,\overline{\mathcal{P}}^{1/2}(\Xi,\epsilon),m):=\int_{\epsilon^{2}/2^{13}}^{\epsilon}\sqrt{H_{B}(u,\overline{\mathcal{P}}^{1/2}(\Xi,\epsilon),m)}du\vee\epsilon,$ where $a\vee b:=\max\{a,b\}$. For brevity, we may omit the dependence on $m$ when it is clear from context.
@@
$\sup_{G_{*}\in\Xi}\mathbb{E}_{p_{G_{*},n}}\mathbb{E}_{X}[d_{H}(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\leq C\sqrt{\log n/n}.$
@@
$H_{B}(\epsilon,\mathcal{P}(\Xi),d_{H})\lesssim\log(1/\epsilon),$ (55)
@@
    $H_{B}(\epsilon,\mathcal{P}(\Xi),d_{H})\lesssim\log(1/\epsilon)$.
@@
Now we will prove Lemma 3, Lemma 4 and Lemma 5 in order. At first we need to introduce another Lemma 6 before we prove Lemma 3. Lemma 6 is Theorem 5.11 in [Vandegeer-2000] and its proof can also be found in [Vandegeer-2000].
@@
Let $R>0$, $k\geq 1$ and $\mathcal{G}$ is a subset in $\Xi$ where $G_{*}\in\mathcal{G}\subset\Xi$ . Given $C_{1}<\infty$, for all $C$ sufficiently large, and for $n\in\mathbb{N}$ and $t>0$ is in the following range
@@
$t\geq C^{2}(C_{1}+1)\Bigg(R\vee\int^{R}_{t/(2^{6}\sqrt{n})}H_{B}^{1/2}\big(\frac{u}{\sqrt{2}},\overline{\mathcal{P}}^{1/2}(\Xi,R),m\big)du\Bigg),$ (57)
@@
Firstly, by Lemma 4.1 and 4.2 in [Vandegeer-2000], we have
$\frac{1}{16}\mathbb{E}_{X}[d_{H}^{2}(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\leq\mathbb{E}_{X}[d_{H}^{2}(\bar{p}_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\leq\frac{1}{\sqrt{n}}\mu_{n}(\widehat{G}_{n}),$
@@
$\mathbb{P}_{G_{*,n}}(\mathbb{E}_{X}[d_{H}(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\geq\delta)$
$\leq\mathbb{P}_{G_{*,n}}\left(\mu_{n}(\widehat{G}_{n})-\sqrt{n}\mathbb{E}_{X}[d_{H}^{2}(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\geq 0,\mathbb{E}_{X}[d_{H}(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\geq\frac{\delta}{4}\right)$
$\leq\mathbb{P}_{G_{*,n}}\left(\sup_{G:\mathbb{E}_{X}[d_{H}(\bar{p}_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]\geq\delta/4}\left[\mu_{n}(G)-\sqrt{n}\mathbb{E}_{X}[d_{H}^{2}(\bar{p}_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]\right]\geq 0\right)$
$\leq\sum_{s=0}^{S}\mathbb{P}_{G_{*,n}}\left(\sup_{G:2^{s}\delta/4\leq\mathbb{E}_{X}[d_{H}(\bar{p}_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]\leq 2^{s+1}\delta/4}\left|\mu_{n}(G)\right|\geq\sqrt{n}2^{2s}(\frac{\delta}{4})^{2}\right)$
$\leq\sum_{s=0}^{S}\mathbb{P}_{G_{*,n}}\left(\sup_{G:\mathbb{E}_{X}[d_{H}(\bar{p}_{G}(\cdot|X),p_{G_{*}}(\cdot|X))]\leq 2^{s+1}\delta/4}\left|\mu_{n}(G)\right|\geq\sqrt{n}2^{2s}(\frac{\delta}{4})^{2}\right)$
where $S$ is a smallest number such that $2^{S}\delta/4>1$ .
@@
$\int^{R}_{t/2^{6}\sqrt{n}}H_{B}^{1/2}\left(\frac{u}{\sqrt{2}},\mathcal{P}^{1/2}(\Xi,R),\mu\right)du\vee 2^{s+1}\delta$
$=\sqrt{2}\int^{R/\sqrt{2}}_{R^{2}/2^{13}}H_{B}^{1/2}\left({u},\mathcal{P}^{1/2}(\Xi,R),\mu\right)du\vee 2^{s+1}\delta$
@@
$\mathbb{P}_{G_{*,n}}\left(\mathbb{E}_{X}[d_{H}(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]>\delta\right)\leq C\sum_{s=0}^{\infty}\exp\left(-\frac{2^{2s}n\delta^{2}}{2^{14}C^{2}}\right)\leq c\exp\left(-\frac{n\delta^{2}}{c}\right),$ (59)
@@
$\mathbb{E}_{p_{G_{*,n}}}\mathbb{E}_{X}[d_{H}({p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X)})]$ $=\int^{\infty}_{0}\mathbb{P}\left(\mathbb{E}_{X}[d_{H}({p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X)})]>\delta\right)d\delta$
$\leq\delta_{n}+c\int^{\infty}_{\delta_{n}}\exp\left(-\frac{n\delta^{2}}{c^{2}}\right)d\delta$
@@
$\sup_{G_{*}\in\Xi}\mathbb{E}_{p_{G_{*},n}}\mathbb{E}_{X}[d_{H}(p_{\widehat{G}_{n}}(\cdot|X),p_{G_{*}}(\cdot|X))]\leq C\sqrt{\log n/n}.$
@@
Now, using the fact that for densities $f^{*},f_{1},f_{2}$, we have $h^{2}\left(\frac{f_{1}+f^{*}}{2},\frac{f_{2}+f^{*}}{2}\right)\leq\frac{h^{2}(f_{1},f_{2})}{2}$, it is easy to verify that $H_{B}(\delta/\sqrt{2},\overline{\mathcal{P}}(\Xi),d_{H})\leq H_{B}(\delta,\mathcal{P}(\Xi),d_{H})$. Hence, if equation (55) holds true, then
$H_{B}(\delta,\overline{\mathcal{P}}^{1/2}(\Xi,\delta),\mu)\leq H_{B}(\delta,{\mathcal{P}}(\Xi),d_{H})\lesssim\log\left(\frac{1}{\delta}\right).$
@@
Let $\mathcal{P}(\Theta):=\{p_{\Upsilon}:\Upsilon\in\Theta\}$, where $p_{\Upsilon}(Y|X):=f(Y|h(X,\eta),\nu)$. By Lemma 6 in [ho2022gaussian], we have
@@
[…]
@@
here $C_{1}$ is a positive constant depending on $\ell$ and $f_{0}$. Moreover $H(X,Y)$ is an envelope of $\mathcal{P}(\Xi)$. Next, let $g_{1},\dots,g_{N}$ represent an $\eta$-net over $\mathcal{P}_{k}(\Xi)$. Then, we construct the brackets $p^{L}_{i}(X,Y),p^{U}_{i}(X,Y)]$ as follows: